Time it took to break a car at 30 mph, 1 second. Calculate the average braking acceleration ¯a (m/s2 ) and record the result (to the nearest 0.1 m/s2 ) as the answer to part (b) of this problem. Also calculate the average braking acceleration in terms of g (g = 9.80 m/s2 ) and record the result (to the nearest 0.01g) as the answer to part (c) of this problem. Note: Enter all of your answers as positive values.
Answer #1

initial velocity = 30mph= 13.4112m/s

final velocity=0 a=v-u/t==0-13.4112/1=13.4112 m/s^2

b)13.4m/s^2

c) 13.4112/9.8 =1.36g